A pilot wants to fly due north a distance of 125 km. The wind is blowing out of the west at a constant 35 km/h. If the plane can travel at 175 km/h, how long will the trip take?
Answer by stanbon(75887)

A pilot wants to fly due north a distance of 125 km. The wind is blowing out of the west at a constant 35 km/h. If the plane can travel at 175 km/h, how long will the trip take?
============
Draw the picture.
You have a right triangle with
leg1 = 125 km
leg2 = 35t
hypotenuse = 175t
-------------------------
(175t)^2 = (35t)^2 + 125^2
29400t^2 = 15625
t^2 = 0.53146...
time = 0.7290 hrs
time = 43.74 minutes
=======================
Cheers,
Stan H.